# Quenched pair breaking by interlayer correlations as a key to superconductivity in La3Ni2O7

### v4

@@ Abstract
We finally discuss implications of the above pressure-induced FS change for the available experimental data. Since the SDW is known to emerge in the ambient-pressure phase [44, 45, 46, 47, 48, 49, 50, 51], direct comparison of our FS for small $|t^{z}_{\perp}|$ with experimental FS obtained from angle-resolved photoemission spectroscopy under ambient pressure [59] may be misleading. Also, considering that there is a discrepancy as to whether or not the $\gamma$ pocket crosses the Fermi level in La4Ni3O10 even between experiments [60, 61], the same issue may also pertain to La3Ni2O7. Further study is required. Rather, a tantalizing signature of the Lifshitz transition of FS is seen in the pressure dependence of Hall coefficient ($R_{\mathrm{H}}$) [62]. While the sign of $R_{\mathrm{H}}$ is positive for the entire pressure range, a sudden drop of $R_{\mathrm{H}}$ occurs near $\sim 10$ GPa followed by the emergence of superconductivity [62]. Since the $\gamma$ pocket is a hole-like FS [Fig. 2(d)] and is destructive for pairing, the drop of $R_{\mathrm{H}}$ and the emergence of superconductivity is quite naturally explained from our Lifshitz transition scenario.
@@
Acknowledgments. We are grateful to F. Lechermann and I. Eremin for useful discussion. S. R. thanks Se Young Park for helpful comment on the charge self-consistency. This work is supported by the Cluster of Excellence ‘CUI: Advanced Imaging of Matter’ of the Deutsche Forschungsgemeinschaft (DFG) - EXC 2056 - project ID 390715994, by DFG priority program SPP 2244 (WE 5342/5-1 project No. 422707584) and the DFG research unit FOR 5242 (WE 5342/7-1, project No. 449119662). Calculations were done on the supercomputer Lise at NHR@ZIB as part of the NHR infrastructure under the project hhp00056.

### v5

@@ Abstract
We finally discuss implications of the above pressure-induced FS change for the available experimental data. Since the SDW is known to emerge in the ambient-pressure phase [44, 45, 46, 47, 48, 49, 50, 51], direct comparison of our FS for small $|t^{z}_{\perp}|$ with experimental FS obtained from angle-resolved photoemission spectroscopy under ambient pressure [59] may be misleading. Also, considering that there is a discrepancy as to whether or not the $\gamma$ pocket crosses the Fermi level in La4Ni3O10 even between experiments [60, 61], the same issue may also pertain to La3Ni2O7. Further study is required. Rather, a tantalizing signature of the Lifshitz transition of FS is seen in the pressure dependence of Hall coefficient ($R_{\mathrm{H}}$) [3, 62]. While the sign of $R_{\mathrm{H}}$ is positive for the entire pressure range, a drop of $R_{\mathrm{H}}$ occurs near $\sim 10$ GPa followed by the emergence of superconductivity [3, 62]. Since the $\gamma$ pocket is a holelike FS [Fig. 2(d)] and is destructive for pairing, the drop of $R_{\mathrm{H}}$ and the emergence of superconductivity is quite naturally explained from our Lifshitz transition scenario.
@@
Acknowledgments. We are grateful to F. Lechermann and I. Eremin for useful discussion. S. R. thanks Se Young Park for helpful comments on the charge self-consistency. This work is supported by the Cluster of Excellence ‘CUI: Advanced Imaging of Matter’ of the Deutsche Forschungsgemeinschaft (DFG) - EXC 2056 - project ID 390715994, by DFG priority program SPP 2244 (WE 5342/5-1 project No. 422707584) and the DFG research unit FOR 5242 (WE 5342/7-1, project No. 449119662). Calculations were done on the supercomputer Lise at NHR@ZIB as part of the NHR infrastructure under the project hhp00056.